1. One inlet pipe can fill an empty pool in 6 hours, and a drain can empty the pool in 8 hours. How long will it take the pipe to fill the pool if the drain is left open?

2. Hannah can paint a room in 8 hours. Destiny can paint the same room in 6 hours. How long does it take for both Hannah and Destiny to paint the room it they are working together.
3.A plane can fly 360 miles in the same time as it takes a car to go 90 miles. If the car travels 90 mph slower than the plane, find the speed of the plane.

Answer by Edwin McCravy(20081)

1. One inlet pipe can fill an empty pool in 6 hours, and a
drain can empty the pool in 8 hours. How long will it take
the pipe to fill the pool if the drain is left open?
               Part of pool   =   Rate in   times  Time in
                filled            pool/hr           hours   
Inlet pipe         1                1/6               6
Outlet pipe       -1               -1/8               8
Both open          1                1/T               T

Rate of Inlet + Rate of outlet = Together rate
       1/6    +   (-1/8)       =  1/T

Multiply by 24T
   
       4T    -       3T        =  24
                      T        =  24

It'll take 24 hours. 

2. Hannah can paint a room in 8 hours. Destiny can paint
the same room in 6 hours. How long does it take for both
Hannah and Destiny to paint the room it they are working
together?
               Part of room       Rate in         Time in
                 painted          room/hr           hours   
Hannah             1                1/6               6
Destiny            1                1/8               8
Both painting      1                1/T               T

Rate of Hannah + Rate of Destiny = Together rate
       1/6    +        1/8       =  1/T

Multiply by 24T
   
       4T    +       3T        =  24
                     7T        =  24
                      T        =  24/7 or 3 3/7 hours.  
It'll take 3 3/7 hours. 

3.A plane can fly 360 miles in the same time as it
takes a car to go 90 miles. If the car travels
90 mph slower than the plane, find the speed of
the plane.
          Distance   =    Rate in       Time in
          traveled       miles/hour     hours = dist./rate   
Plane        360              x           360/x
Car           90             x-90        90/(x-90)


          Plane's time = Car's time.

                 360/x = 90/(x-90)

Cross multiply

             360(x-90) = 90x
            360x-32400 = 90x
                  270x = 32400
                     x = 32400/270
                     x = 120 miles/hour

Edwin